# MAGICS I. The First Few Orbits Encode the Fate of Seed Massive Black Hole Pairs

### v1

@@ 5 Results
Using the method described in Section 4, we set up a total of 15 galaxies and MBH merger initial conditions for the chosen ASTRID merging systems shown in Section 3, and with the “full-physics” subgrid physics models depicted in Section 2. In this section, we show the results from these high-resolution resimulations.
@@ 5.1 Evolution of the host galaxies
We summarize the evolution in central stellar density across all resimulated systems in Figure 11, from the start of the resimulation ($\sim 300\,{\rm Myrs}$ before the ASTRID merger) to the resimulation merger time. The thick colored lines show the systems that merged in the resimulation, while the thin grey lines show the systems that stall at $\sim kpc$ scales for more than $1.5\,{\rm Gyr}$. For almost all systems the stellar densities increase by an order of magnitude during the MBH sinking. After the ASTRID merger, the stellar density still increases by a factor of $2\sim 3$. This direct comparison implies that one should account for the newly formed stars when using post-processed analytical models to compute the binary hardening time for mergers in cosmological simulations, and the resulting hardening efficiency may increase. We also note that for the system with the steepest increase in central stellar density (system6, shown in thick grey), the MBH sinks even more efficiently in the resimulation system and merges before the ASTRID merger takes place.
The recent work by Liao et al. (2023b) has shown that the increase in central stellar density and the development of a nuclear stellar core can significantly increase the binary hardening efficiency in SMBHs. Our result again highlights the effect of central star formation on MBH sinking timescales, in the context of high-redshift seed-mass MBH mergers. In the resimulation runs, we include the “full-physics” modeling of both star formation and AGN feedbacks (thermal and kinetic), and thus the stellar density profile is more “core-like” (similar to the “CoolStarKinAGN” modeling in Liao et al. (2023b)). Nonetheless, we still find very high star formation rates (up to $50\,M_{\odot}/{\rm yr}$) and the development of stellar nuclei when the merging galaxies are compact enough (e.g. system11), and the MBH seeds can sink very efficiently in these cases.
Figure 11: Top panel: Evolution of the central stellar density (measured at $100\,{\rm pc}$ from the most massive galaxy center) during the MBH inspiral and merger in the simulations. The crosses mark the simulation MBH merging time, and the vertical dotted line marks the ASTRID merger time. Bottom panel: the ratio between stellar density measured at the resimulation merger and the initial condition (green), and between the resimulation merger and the ASTRID merger (orange).
@@ 5.2 Seed MBH merging timescale
In ASTRID, MBH orbits are resolved down to scales of $\sim 1\,{\rm kpc}$, and so MBH pairs are assumed to “merge” after that. However, in high-redshift dwarf galaxies, it is typical for seed MBHs to stall on kpc scale orbits for over a few Gyrs (e.g. Ma et al., 2021; Pfister et al., 2019; Partmann et al., 2023). In this section, we study the merging timescales and the stalling of MBH seeds of the systems in the resimulation, and the correlation with large-scale orbital and galaxy properties.
@@
We note that in the resimulation we still do not resolve the full dynamical range until the MBH coalescence, and so the MBHs “merge” when their orbital separation is $\sim 20\,{\rm pc}$. Nonetheless, according to previous works that follow the MBH orbital evolution to the hardening phase, MBH pairs generally enter into the binary hardening regime very rapidly after they sink below $<20\,{\rm pc}$ (e.g. Gualandris et al., 2022; Partmann et al., 2023; Koehn et al., 2023). Further stalling may happen at $<pc$ scales due to the depletion of the loss cone, but here we only focus on the pair evolution in the dynamical friction regime and defer the smaller scale dynamics to future works.
Figure 13: Left panel: relation between MBH merging timescale and stellar density at $50\,{\rm pc}$ from the galaxy center. We measure density both at the ASTRID merger time (green) and at the resimulation merger time (pink). We see a tight correlation between the merging timescale and stellar densities. We plot densities in systems that do not merge within 1.5 Gyrs in the resimulation on the right of the box. Right panel: initial (green) and final (pink) orbital eccentricity of MBH pairs in the resimulation and the correlation with the merging timescale. The eccentricity cluster around 0.8 when the pair start entering into the hardening phase.
@@ 5.3 MBH mergers and large-scale properties
One main motivation for using cosmological simulations to set up resimulations of galaxy and MBH mergers is to understand whether we can use the information from cosmological simulations to predict the dynamics of MBHs at sub-resolution scales. In this section, we connect each resimulation and the MBH merging time back to properties of the ASTRID-resolved quantities, and investigate what would be a good indicator for the sub-resolution dynamical behavior of MBH seeds.
@@
Figure 17: Left column: Rate of total energy change of the secondary MBH (pink) compared with the energy loss rate due to dynamic friction (blue). The lines show each system and the diamonds are the median across all systems in each time bin. We plot the systems that merged in the resimulation in the top panel and the systems that stalled in the bottom panel. The merged MBHs experience loss of energy dominated by gravity, while the stalled MBHs gain energy during the first $\sim 200\,{\rm Myrs}$ of the galaxy merger. Middle column: Cumulative change in the MBH energy since the first pericentric passage between the pair. Right column: visualization of two galaxy mergers that lead to a merged pair (top) and a stalled pair (bottom). The stalled MBHs are mostly found in head-on collisions of minor galaxy mergers, in which the secondary host is quickly dissolved.
@@
Finally, Figure 17 shows the change in the total energy of the secondary MBH after the first pericentric passage between the pairs, compared with the energy loss due to dynamical friction (in the resimulation). We find a striking contrast between the merged and stalled system at the very early stage of the merger: the merged MBHs lost most of their energy within the first $\sim 100\,{\rm years}$. The energy loss due to dynamical friction then begins to take effect after $\sim 200\,{\rm Myrs}$ to further drive the merger. In contrast, the stalled MBHs gained energy in this beginning phase from the gravitational torque. As a result, the MBHs never made their way into the central region where dynamical friction acts effectively. On the right, we show a typical example of both the merged and stalled scenarios. Consistent with the picture in Figure 14, the MBHs in major mergers experience energy loss at the potential center of their hosts, whereas those in minor mergers with head-on collisions gain energy from the tidal disruption of their host galaxies. In future development of subgrid merger models in cosmological simulations, it will be useful to measure the energy change rates of MBHs as an indication of the merging timescales and the likelihood of stalling.

### v3

@@ 5 Results
Using the method described in Section 4, we set up a total of 15 galaxies and MBH merger initial conditions for the chosen ASTRID merging systems shown in Section 3, and with the “full-physics" subgrid physics models depicted in Section 2. In this section, we show the results from these high-resolution resimulations.
@@ 5.1 Evolution of the host galaxies
We summarize the evolution in central stellar density across all resimulated systems in Figure 11, from the start of the resimulation ($\sim 300\,{\rm Myrs}$ before the ASTRID merger) to the resimulation merger time. The thick colored lines show the systems that merged in the resimulation, while the thin grey lines show the systems that stall at $\sim kpc$ scales for more than $1.5\,{\rm Gyr}$. For almost all systems the stellar densities increase by an order of magnitude during the MBH sinking. After the ASTRID merger, the stellar density still increases by a factor of $2\sim 3$. This direct comparison implies that one should account for the newly formed stars when using post-processed analytical models to compute the binary hardening time for mergers in cosmological simulations, and the resulting hardening efficiency may increase. We also note that for the system with the steepest increase in central stellar density (system11, shown in light brown), the MBH sinks even more efficiently in the resimulation system and merges before the ASTRID merger takes place.
The recent work by Liao et al. (2023b) has shown that the increase in central stellar density and the development of a nuclear stellar core can significantly increase the binary hardening efficiency in SMBHs. Our result again highlights the effect of central star formation on MBH sinking timescales, in the context of high-redshift seed-mass MBH mergers. In the resimulation runs, we include the “full-physics" modeling of both star formation and AGN feedbacks (thermal and kinetic), and thus the stellar density profile flattens (similar to the “CoolStarKinAGN" modeling in Liao et al. (2023b)). Nonetheless, we still find very high star formation rates (up to $50\,M_{\odot}/{\rm yr}$) and the development of stellar nuclei when the merging galaxies are compact enough (e.g. system11), and the MBH seeds can sink very efficiently in these cases.
Figure 11: Top panel: Evolution of the central stellar density (measured at $100\,{\rm pc}$ from the most massive galaxy center) during the MBH inspiral and merger in the simulations.The colored lines show the density evolution until the resimulation merger time (crosses) for systems that merged in the resimulations. The thin grey lines are systems that stall in the resimulations. The vertical dotted line marks the ASTRID merger time. Bottom panel: the ratio between stellar density measured at the resimulation merger and the initial condition (green), and between the resimulation merger and the ASTRID merger (orange).
@@ 5.2 Seed MBH merging timescale
In ASTRID, MBH orbits are resolved down to scales of $\sim 1\,{\rm kpc}$, and so MBH pairs are assumed to “merge" after that. However, in high-redshift dwarf galaxies, it is typical for seed MBHs to stall on kpc scale orbits for over a few Gyrs (e.g. Ma et al., 2021; Pfister et al., 2019; Partmann et al., 2023). In this section, we study the merging timescales and the stalling of MBH seeds of the systems in the resimulation, and the correlation with large-scale orbital and galaxy properties.
@@
We note that in the resimulation we still do not resolve the full dynamical range until the MBH coalescence, and so the MBHs “merge" when their orbital separation is $\sim 20\,{\rm pc}$. To validate that the MBHs’ motion is not affected by numerical noise above the merging distance, we measure the wandering radius of the merger remnant following Bortolas et al. (2016), by averaging the mean displacement of the merger remnant from the galaxy center over time. We find that for all systems that merged, the remnant MBH has a mean displacement of $\sim 20\,{\rm pc}$. By numerically merging the MBH at this separation, we pick out systems that will likely form a bound binary, since the stalling in the dynamical friction regime is seen at $\sim{\rm kpc}$ scales (e.g. Gualandris et al., 2022; Partmann et al., 2023; Koehn et al., 2023). Further stalling may happen at $\sim pc$ scales due to the depletion of the loss cone, but here we only focus on the pair evolution in the dynamical friction regime and defer the smaller scale dynamics to future works.
Figure 13: Left panel: relation between MBH merging timescale and stellar density at $50\,{\rm pc}$ from the galaxy center. We measure density both at the ASTRID merger time (green) and at the resimulation merger time (pink). We see a tight correlation between the merging timescale and stellar densities. We plot densities in systems that do not merge within 1.5 Gyrs in the resimulation on the right of the box. Right panel: initial (green) and final (pink) orbital eccentricity of MBH pairs in the resimulation and the correlation with the merging timescale. The eccentricity clusters around 0.8 when the pair starts entering into the hardening phase.
@@ 5.3 MBH mergers and large-scale properties
One main motivation for using cosmological simulations to set up resimulations of galaxy and MBH mergers is to understand whether we can use the information from cosmological simulations to predict the dynamics of MBHs at sub-resolution scales. In this section, we connect each resimulation and the MBH merging time back to properties of the ASTRID-resolved quantities and investigate what would be a good indicator for the sub-resolution dynamical behavior of MBH seeds.
@@
Figure 17: Left column: Rate of total energy change of the secondary MBH (blue) compared with the energy loss rate due to dynamic friction (pink). The lines show each system and the diamonds are the median across all systems in each time bin. We plot the systems that merged in the resimulation in the top panel and the systems that stalled in the bottom panel. The merged MBHs experience loss of energy dominated by gravity, while the stalled MBHs gain energy during the first $\sim 200\,{\rm Myrs}$ of the galaxy merger. Middle column: Cumulative change in the MBH energy since the first pericentric passage between the pair. Right column: visualization of two galaxy mergers that lead to a merged pair (top) and a stalled pair (bottom). The stalled MBHs are mostly found in head-on collisions of minor galaxy mergers, in which the secondary host is quickly dissolved.
@@
Finally, Figure 17 shows the change in the total energy of the secondary MBH after the first pericentric passage between the pairs, compared with the energy loss due to dynamical friction (in the resimulation). We find a striking contrast between the merged and stalled system at the very early stage of the merger: the merged MBHs lost most of their energy within the first $\sim 100\,{\rm Myrs}$. The energy loss due to dynamical friction then begins to take effect after $\sim 200\,{\rm Myrs}$ to further drive the merger. In contrast, the stalled MBHs gained energy in this beginning phase from the gravitational torque. As a result, the MBHs never made their way into the central region where dynamical friction acts effectively. On the right, we show a typical example of both the merged and stalled scenarios. Consistent with the picture in Figure 14, the MBHs in major mergers experience energy loss at the potential center of their hosts, whereas those in minor mergers with head-on collisions gain energy from the tidal disruption of their host galaxies. In future development of subgrid merger models in cosmological simulations, it will be useful to measure the energy change rates of MBHs as an indication of the merging timescales and the likelihood of stalling.